OpenStudy (anonymous):

Adam is approved for a 25-year 6/1 ARM at 4.35% with a 4/14 cap structure in the amount of $308,400. Assuming Adam is charged the maximum interest rate, what will his monthly payments be at the beginning of the seventh year? $1,431.85 $1,688.04 $2,080.06 $2,291.86

OpenStudy (anonymous):

I think first I gotta find the maximum interest rate on the loan .

OpenStudy (anonymous):

4.35 % + the cap structure which is 14

OpenStudy (amistre64):

the interest is held fixed for 6 years, then goes up 4 points each year (not to exceed 14 points in all)

OpenStudy (amistre64):

determine the payments for the first 6 years to define the balance after 6 years then restructure the loan for the remaining time at +4 points interest

OpenStudy (amistre64):

payments for a 25yr, fixed loan of .0435 will give us the initial period payments amounts Po = 308400k^(12*25)(1-k)/(1-k^(12*25)), k=1+.0435/12 B6 = 308400k^(12*6)-Po(1-k^(12*6))/(1-k), k=1+.0435/12 use that for the remaining time (25-6) years at .0835 P1 = B6 k^(12*19)(1-k^(12*19))\(1-k), k=1+.0835/12

OpenStudy (amistre64):

Po = 1688.04 B6 = 261 595 i see i typoed P1; that last fraction is flipped the wring way :) P1 = 261595 k^(12*19)(1-k)/(1-k^(12*19)), k=1+.0835/12 = 2292 or so

OpenStudy (anonymous):

@amistre64 Thank you alot